y''+3y'-2y=e-x/x


1
Expert's answer
2021-08-23T17:51:10-0400

"y'' + 3y - 2y = \frac{ e^{-x}}{x}.\\
\text{This is a second order non-homogeneous DE}\\
\text{The characteristics equation is ;}\\
m^2+3m-2=0\\
m=\frac{-3 \pm \sqrt{3^2-4(1)(-2)}}{2(1)}\\
m=\frac{-3 \pm \sqrt{17}}{2}\\

\text{The complimentary solution is;}\\
y_c= Ae^{\frac{-3 + \sqrt{17}}{2}x} + Be^{\frac{-3 - \sqrt{17}}{2}x}.\\
\text{To get he Particular solution is;}\\
y_p=C_1y_1+C_2y_2\\
\text{Where:} ~~~~C_1=\int \frac{-y_2f(x)}{W(x)} ~~ \text{and}~~ C_2=\int \frac{y_1f(x)}{W(x)}\\
W(x)=y_1y_2'-y_1'y_2\\
W(x)=\left(\frac{-3 - \sqrt{17}}{2}x\right)e^{\frac{-3 + \sqrt{17}}{2}x}e^{\frac{-3 - \sqrt{17}}{2}x}-\left(\frac{-3 + \sqrt{17}}{2}x\right)e^{\frac{-3 + \sqrt{17}}{2}x}e^{\frac{-3 - \sqrt{17}}{2}x}=-\sqrt{17} e^{-3x}\\\\
f(x)=\frac{e^{-x}}{x}\\\\
\\C_1=\int\frac{e^{\frac{-3 + \sqrt{17}}{2}x}}{x\sqrt{17}}dx=\frac{\operatorname{Ei}{\left(\frac{ 1 - \sqrt{17}}{2}x \right)}}{\sqrt{17}}+C\\
C_2=\int\frac{e^{\frac{7 - \sqrt{17}}{2}x}}{-x\sqrt{17}}dx=-\frac{\operatorname{Ei}{\left(\frac{ 7- \sqrt{17}}{2}x \right)}}{\sqrt{17}}+C\\\\
\text{The general solution is given as;}\\
y(x)=Ay_1+By_2+C_1y_1+C_2y_2\\\\
y(x)=Ae^{\frac{-3 + \sqrt{17}}{2}x} + Be^{\frac{-3 - \sqrt{17}}{2}x}+\frac{\operatorname{Ei}{\left(\frac{ 1 - \sqrt{17}}{2}x \right)}}{\sqrt{17}}e^{\frac{-3 + \sqrt{17}}{2}x}-\frac{\operatorname{Ei}{\left(\frac{ 7- \sqrt{17}}{2}x \right)}}{\sqrt{17}}e^{\frac{-3 - \sqrt{17}}{2}x}."